# Collaborative Representation Learning for Alignment of Tactile, Language, and Vision Modalities

### v1

@@ 1 Introduction
Tactile is one of the essential senses of human perception. Through tactile interaction, we can sense both static and dynamic attributes of objects (e.g., material texture, roughness, and hardness), many of which are too subtle to be reliably perceived by other perception systems like vision (cheng2024touch100k; shi2025presentagent; dave2024multimodal; li2025madakv; li2025flowmm; jiang2025arnet). In recent years, researchers have been striving to help robots understand the complex and realistic physical world by designing high-resolution tactile sensors (yuan2017gelsight; donlon2018gelslim; lambeta2020digit; gelsight; zhang2024duragel) comparable to human touch and collecting large-scale indoor and outdoor tactile image datasets (yang2022touch; kerr2022ssvtp; fu2024tvl; yu2024octopi; tacquad).
@@
However, tactile sensors are not yet fully standardized. Due to external factors (e.g., camera type, lighting position, color, and illumination), tactile images can differ significantly even under identical touch object conditions (Fig. 1(i) and (ii)). To address these variations, previous studies (yang2024binding; tacquad) have borrowed the concept of positional encoding from language models (su2024roformer; zhao2023length; jiang2025purekv), introducing learnable tokens to model sensor-specific characteristics. However, these methods overlook a crucial fact: even when touch objects differ noticeably, the styles of the tactile images can still be quite similar (Fig. 1(iii)), which poses a challenge for tactile models to disentangle tactile-irrelevant features.
@@
Vision and language are also core channels for human-environment interaction. In real-world tasks, their integration has been extensively studied. A large body of work (radford2021learning; jia2021scaling; alayrac2022flamingo; li2023scaling) has successfully built semantic bridges between visual and linguistic modalities through contrastive learning (oord2018representation), achieving remarkable progress. This success has since extended to additional modalities, including audio, point clouds, event etc (girdhar2023imagebind; guo2023point; wang2024omnibind; lyu2024unibind). Despite the flourishing development of multimodal learning catalyzed by vision-language pretraining, the tactile modality remains significantly underexplored.
@@
Recent research on tactile-language-vision learning has focused on representation learning based on CLIP (radford2021learning; cherti2023reproducible). For instance, TLV-Link (cheng2024touch100k), designed specifically for the GelSight sensor (yuan2017gelsight), trains a tactile encoder via curriculum learning to achieve effective tri-modal alignment. AnyTouch (tacquad) proposes a unified representation learning framework for static-dynamic and multi-sensor tactile data, employing masked modeling, self-supervised multimodal alignment, and cross-sensor matching to improve generalization across different sensors. Vit-Lens-2 (lei2024vitlen) introduces a generic multimodal encoding approach that first transforms various modality inputs into intermediate representations using lightweight modules and then feeds them into a frozen pretrained ViT (dosovitskiy2020image), enabling efficient representation learning. However, these methods face two challenges: (1) Most methods adapt LoRA (hu2022lora; zhou2025cola) within single modality branches, without explicitly modeling the synergy among the three modalities before fusion, limiting their deep fusion capability. (2) There is a lack of standardized evaluation settings (e.g., base models, batch sizes), making fair comparisons difficult.
@@
Figure 1: Three properties of heterogeneous sensors are identified: (i). Tactile sensors lack full standardization, leading to significant tactile images variation (yang2024binding). (ii). Tactile images from the identical touch object can differ inconsistently (e.g., ⓐ and ⓑ are similar, both differing greatly from ⓒ). (iii). Despite different touch objects, tactile images may share a consistent style (e.g., ⓐ and ⓑ resemble ⓓ in a dark tone tinged with red).
@@
To this end, we propose a CLIP-based Tactile-Language-Vision Collaborative Representation (TLV-CoRe) learning method, aiming to learn sensor-agnostic tactile representations. Unlike existing methods (yang2024binding; tacquad), TLV-CoRe introduces a learnable Sensor-Aware Modulator (SAM) that adaptively maps tactile features from different sensors into a unified parameter space, combined with tactile-irrelevant decoupled learning to effectively disentangle tactile-irrelevant features. Furthermore, to strengthen cross-modal collaboration in intermediate representations, TLV-CoRe incorporates a Unified Bridging Adapter (UBA) into three modality encoders. UBA consists of modality-specific projection layers to learn individual representations and a shared feature space mapping layer to facilitate tri-modal alignment.
@@
For evaluation, we propose a RSS evaluation framework, designed to analyze the Robustness, Synergy, and Stability of various tactile representation learning methods. We define three evaluation protocols—intra-sensor evaluation, cross-sensor generalization, and multi-sensor generalization—to assess the robustness of different methods. Given that multimodal alignment should enhance rather than compromise individual modality performance (wang2022deep; jiang2025fedcfa; li2025mergenet; dufumier2024align), we introduce modal cross-evaluagtion tasks (especially between tactile and vision modalities) to assess synergy of various modal encoders. We also investigate the impact of batch size on model stability, as varying batch sizes affect the number of negative samples in contrastive learning. We encourage future research to adopt the RSS framework for comprehensive comparisons of different multimodal tactile methods based on CLIP.
@@ 2 Related Works
TLV-Link (cheng2024touch100k) CLIP-Based
AnyTouch (tacquad) CLIP-Based
VIT-LENS-2 (lei2024vitlen) CLIP-Based
UniTouch (yang2024binding) CLIP-Based / LLM-Based
TVL-LLaMA (fu2024tvl) LLM-Based
VT CMC (yang2022touch) Custom-Based
T3 (zhao2024transferable) Custom-Based
MViTac (dave2024multimodal) Custom-Based
TLA (hao2025tla) Custom-Based
SITR (gupta2025sensor) Custom-Based
@@ 2.1 Multimodal Alignment
Multimodal alignment (baltruvsaitis2018multimodal; xu2023multimodal) aims to build bridges between different modalities. It not only helps models better understand cross-modal information, but indirectly facilitates representation learning within individual modalities (zhou2025revisiting; zhou2025disentangled; zhou2025cuff; zhang2024vision). The success of CLIP (radford2021learning) has sparked rapid development in vision-language pretraining methods (jia2021scaling; alayrac2022flamingo; li2023scaling; zhao2023mmicl). Subsequent research has extended alignment to additional modalities, e.g., audio (guzhov2022audioclip), video (ma2022x), and 3D point clouds (xue2023ulip). ImageBind (girdhar2023imagebind) constructs a unified embedding space across six modalities through image-pairing learning, achieving impressive results in both visual and non-visual tasks. Inspired by ImageBind, Point-Bind (guo2023point) aligns 3D point clouds with 2D images, text, audio, and video by constructing a joint embedding space. UniBind (lyu2024unibind) further aligns multimodal embeddings to a large-model-enhanced embedding center via contrastive learning, resulting in a unified and balanced representation space. Most of these methods adopt the InfoNCE contrastive learning paradigm (oord2018representation), and have demonstrated significant performance improvements, underscoring the effectiveness of this approach. Following this direction, we extend the concept of multimodal alignment to the relatively underexplored tactile modality, leveraging CLIP’s powerful vision-language pretraining capabilities.
@@ 2.2 Tactile Representation Learning
In recent years, the large-scale collection of tactile datasets (yang2022touch; kerr2022ssvtp; fu2024tvl; yu2024octopi; cheng2024touch100k; tacquad) and advances in tactile sensors (yuan2017gelsight; lambeta2020digit; gelsight; zhang2024duragel) have significantly accelerated research in tactile representation learning. VT CMC (yang2022touch) models tactile images from GelSight sensors (yuan2017gelsight) using contrastive multiview coding (tian2020contrastive). T3 (zhao2024transferable) proposes an architecture with sensor-specific encoders, a shared backbone network, and task-specific decoders, enabling transferable tactile representation learning across multiple sensors. MViTac (dave2024multimodal) uses an InfoNCE loss to jointly optimize visual and tactile features for effective intra- and inter-modal fusion. TLA (hao2025tla) encodes temporal tactile images into composite images and incorporates language reasoning for cross-modal finetuning, thereby promoting generalized tactile-language-action policy learning. SITR (gupta2025sensor) combines supervised contrastive learning with physics-based simulation to learn sensor-invariant representations, enabling zero-shot transfer across GelSight sensors. UniTouch (yang2024binding) aligns tactile signals with visual data and introduces learnable sensor-specific tokens to leverage CLIP for multimodal shared representations, supporting various zero-shot tactile tasks. Additionally, UniTouch integrates large language models to facilitate diverse tactile question-answering tasks. TVL-LLaMA (fu2024tvl) proposes a tri-modal contrastive-trained tactile encoder aligned with vision and language, and further finetunes LLaMA2 to generate tactile descriptions from visual and tactile inputs. TLV-Link (cheng2024touch100k) combines teacher-student curriculum learning with contrastive learning for tactile-centric multimodal pretraining. AnyTouch (tacquad) integrates static and dynamic information through a hierarchical architecture, incorporating masked modeling, multimodal alignment, and cross-sensor matching for unified multi-sensor tactile representation learning. VIT-LENS-2 (lei2024vitlen) leverages a pretrained ViT (dosovitskiy2020image) and modality-specific lens modules for efficient, scalable multimodal tactile learning.
As shown in Table 1, the aforementioned methods fall into three categories: CLIP-Based, LLM-Based, and Custom-Based. The trend shows a shift towards custom end-to-end architectures, but this makes fair benchmarking difficult. Moreover, variations in batch size can significantly affect the stability of multimodal models on tactile tasks (higuera2024sparsh). Our work focuses on CLIP-Based approaches, which allows us to use a consistent base model (cherti2023reproducible) and build a standardized RSS evaluation framework for fairer comparisons.
@@ 3 Methodology
We introduce TLV-CoRe, a method for learning collaborative representations that align the tactile (T), visual (V), and language (L) modalities in a shared latent space (see Fig. 2). TLV-CoRe comprises separate encoders for each modality and two modules: a Sensor-Aware Modulator (SAM) in the tactile branch, and a cross-modal Unified Bridging Adapter (UBA) that aligns the modalities. The SAM removes sensor-specific biases (zhao2024transferable; yang2024binding) from tactile features, yielding invariant representations across various tactile sensors. UBA is a lightweight module that projects features from each modality into a common latent space to facilitate alignment. We train the model with symmetric contrastive losses between modalities together with a sensor-invariance loss to enforce consistency across sensors. Finally, we provide theoretical guarantees of TLV-CoRe’s three properties. Detailed proofs can be found in the Appendix. We describe each component in detail below.
@@ 3.1 Tactile Encoder and Sensor-Aware Modulator
The tactile encoder $\mathcal{E}_{T}$ processes raw tactile inputs $x^{T}$ and produces a $d$-dimensional latent representation. Specifically, we implement $\mathcal{E}_{T}$ as a Vision Transformer (ViT) (dosovitskiy2020image). The tactile image is divided into non-overlapping patches, each patch is linearly embedded, and the resulting sequence is processed through standard transformer blocks. This design enables the encoder to capture spatial patterns and tactile features.
@@
However, tactile images captured under similar conditions can exhibit similar patterns across different sensors (ou2024marker). In such cases, the SAM inadvertently clusters features by sensor identity rather than by underlying tactile content. To address this, we employ tactile-irrelevant decoupled learning that explicitly removes redundant information from the tactile representation.
@@ 3.2 Unified Bridging Adapter for Cross-Modal Alignment
In our implementation we attach exactly L=12L\!=\!12 UBA blocks to each modality so that every sample—regardless of its source—passes through the same number of shared transformations before the final projection. Concretely, the language encoders comprise $12$ transformer layers; we attach one UBA per layer. The vision/tactile branch is deeper (CLIP design (cherti2023reproducible) with $24$ layers), so we leave the first $12$ layers un‐bridged to preserve low‑level visual/tactile primitives and attach UBAs only to the upper $12$ layers where high‑level semantics emerge.
@@ 3.3 Cross-Modal Contrastive Learning
We train TLV-CoRe with symmetric contrastive losses for each pair of modalities and the sensor-invariance loss. For each pair $(X,Y)\in\{(\mathrm{T},\mathrm{V}),(\mathrm{T},\mathrm{L}),(\mathrm{V},\mathrm{L})\}$, we use a symmetric InfoNCE contrastive loss (oord2018representation). Given a batch of $N$ aligned triplets $\{(z^{T}_{i},z^{V}_{i},z^{L}_{i})\}_{i=1}^{N}$, the loss for the tactile–vision pair (as an example) is:

### v3

@@ 1 Introduction
Tactile is one of the essential senses of human perception. Through tactile interaction, we can sense both static and dynamic attributes of objects (e.g., material texture, roughness, and hardness), many of which are too subtle to be reliably perceived by other perception systems like vision (Cheng et al. 2025; Shi et al. 2025; Dave, Lygerakis, and Rueckert 2024; Li et al. 2025a, b; Lv et al. 2025a; Jiang et al. 2025b, a). In recent years, researchers have been striving to help robots understand the complex and realistic physical world by designing high-resolution tactile sensors (Yuan, Dong, and Adelson 2017; Donlon et al. 2018; Lambeta et al. 2020; Zhang et al. 2024b) comparable to human touch and collecting large-scale indoor and outdoor tactile image datasets (Yang et al. 2022; Kerr et al. 2022; Fu et al. 2024; Yu et al. 2024; Feng et al. 2025).
@@
However, tactile sensors are not yet fully standardized. Due to external factors (e.g., camera type, lighting position, color, and illumination), tactile images can differ significantly even under identical touch object conditions (Fig. 1(i) and (ii)). To address these variations, previous studies (Yang et al. 2024; Feng et al. 2025) have borrowed the concept of positional encoding from language models (Su et al. 2024; Zhao et al. 2024b; Jiang et al. 2025c), introducing learnable tokens to model sensor-specific characteristics. However, these methods overlook a crucial fact: even when touch objects differ noticeably, the styles of the tactile images can still be quite similar (Fig. 1(iii)), which poses a challenge for tactile models to disentangle tactile-irrelevant features.
@@
Vision and language are also core channels for human-environment interaction. In real-world tasks, their integration has been extensively studied. A large body of work (Radford et al. 2021; Lv et al. 2025b; Jia et al. 2021; Alayrac et al. 2022; Li et al. 2023) has successfully built semantic bridges between visual and linguistic modalities through contrastive learning (Oord, Li, and Vinyals 2018), achieving remarkable progress. This success has since extended to additional modalities, including audio, point clouds, event etc (Girdhar et al. 2023; Guo et al. 2023; Wang et al. 2024; Lyu et al. 2024). Despite the flourishing development of multimodal learning catalyzed by vision-language pretraining, the tactile modality remains significantly underexplored.
@@
Recent research on tactile-language-vision learning has focused on representation learning based on CLIP (Radford et al. 2021; Cherti et al. 2023). For instance, TLV-Link (Cheng et al. 2025), designed specifically for the GelSight sensor (Yuan, Dong, and Adelson 2017), trains a tactile encoder via curriculum learning to achieve effective tri-modal alignment. AnyTouch (Feng et al. 2025) proposes a unified representation learning framework for static-dynamic and multi-sensor tactile data, employing masked modeling, self-supervised multimodal alignment, and cross-sensor matching to improve generalization across different sensors. Vit-Lens-2 (Lei et al. 2024) introduces a generic multimodal encoding approach that first transforms various modality inputs into intermediate representations using lightweight modules and then feeds them into a frozen pretrained ViT (Dosovitskiy 2020), enabling efficient representation learning. However, these methods face two challenges: (1) Most methods adapt LoRA (Hu et al. 2022; Zhou, Yao, and Chen 2025) within single modality branches, without explicitly modeling the synergy among the three modalities before fusion, limiting their deep fusion capability. (2) There is a lack of standardized evaluation settings (e.g., base models, batch sizes), making fair comparisons difficult.
@@
Figure 1: Three properties of heterogeneous sensors are identified: (i). Tactile sensors lack full standardization, leading to significant tactile images variation (Yang et al. 2024). (ii). Tactile images from the identical touch object can differ inconsistently (e.g., ⓐ and ⓑ are similar, both differing greatly from ⓒ). (iii). Despite different touch objects, tactile images may share a consistent style (e.g., ⓐ and ⓑ resemble ⓓ in a dark tone tinged with red).
@@
To this end, we propose a CLIP-based Tactile-Language-Vision Collaborative Representation (TLV-CoRe) learning method, aiming to learn sensor-agnostic tactile representations. Unlike existing methods (Yang et al. 2024; Feng et al. 2025), TLV-CoRe introduces a learnable Sensor-Aware Modulator (SAM) that adaptively maps tactile features from different sensors into a unified parameter space, combined with tactile-irrelevant decoupled learning to effectively disentangle tactile-irrelevant features. Furthermore, to strengthen cross-modal collaboration in intermediate representations, TLV-CoRe incorporates a Unified Bridging Adapter (UBA) into three modality encoders. UBA consists of modality-specific projection layers to learn individual representations and a shared feature space mapping layer to facilitate tri-modal alignment.
@@
For evaluation, we propose a RSS evaluation framework, designed to analyze the Robustness, Synergy, and Stability of various tactile representation learning methods. We define three evaluation protocols—intra-sensor evaluation, cross-sensor generalization, and multi-sensor generalization—to assess the robustness of different methods. Given that multimodal alignment should enhance rather than compromise individual modality performance (Wang et al. 2022; Jiang et al. 2025d; Wu et al. 2024; Li et al. 2025c; Dufumier et al. 2024), we introduce modal cross-evaluagtion tasks (especially between tactile and vision modalities) to assess synergy of various modal encoders. We also investigate the impact of batch size on model stability, as varying batch sizes affect the number of negative samples in contrastive learning. We encourage future research to adopt the RSS framework for comprehensive comparisons of different multimodal tactile methods based on CLIP.
@@ 2 Related Works
TLV-Link (Cheng et al. 2025) CLIP-Based
AnyTouch (Feng et al. 2025) CLIP-Based
VIT-LENS-2 (Lei et al. 2024) CLIP-Based
UniTouch (Yang et al. 2024) CLIP-Based / LLM-Based
TVL-LLaMA (Fu et al. 2024) LLM-Based
VT CMC (Yang et al. 2022) Custom-Based
T3 (Zhao et al. 2024a) Custom-Based
MViTac (Dave, Lygerakis, and Rueckert 2024) Custom-Based
TLA (Hao et al. 2025) Custom-Based
SITR (Gupta et al. 2025) Custom-Based
@@ 2.1 Multimodal Alignment
Multimodal alignment (Baltrušaitis, Ahuja, and Morency 2018; Xu, Zhu, and Clifton 2023) aims to build bridges between different modalities. It not only helps models better understand cross-modal information, but indirectly facilitates representation learning within individual modalities (Zhou, Han, and Chen 2025; Zhou et al. 2025b, a; Zhang et al. 2024a). The success of CLIP (Radford et al. 2021) has sparked rapid development in vision-language pretraining methods (Jia et al. 2021; Alayrac et al. 2022; Li et al. 2023; Zhao et al. 2023). Subsequent research has extended alignment to additional modalities, e.g., audio (Guzhov et al. 2022), video (Ma et al. 2022), and 3D point clouds (Xue et al. 2023). ImageBind (Girdhar et al. 2023) constructs a unified embedding space across six modalities through image-pairing learning, achieving impressive results in both visual and non-visual tasks. Inspired by ImageBind, Point-Bind (Guo et al. 2023) aligns 3D point clouds with 2D images, text, audio, and video by constructing a joint embedding space. UniBind (Lyu et al. 2024) further aligns multimodal embeddings to a large-model-enhanced embedding center via contrastive learning, resulting in a unified and balanced representation space. Most of these methods adopt the InfoNCE contrastive learning paradigm (Oord, Li, and Vinyals 2018), and have demonstrated significant performance improvements, underscoring the effectiveness of this approach. Following this direction, we extend the concept of multimodal alignment to the relatively underexplored tactile modality, leveraging CLIP’s powerful vision-language pretraining capabilities.
@@ 2.2 Tactile Representation Learning
In recent years, the large-scale collection of tactile datasets (Yang et al. 2022; Kerr et al. 2022; Fu et al. 2024; Yu et al. 2024; Cheng et al. 2025; Feng et al. 2025) and advances in tactile sensors (Yuan, Dong, and Adelson 2017; Lambeta et al. 2020; Zhang et al. 2024b) have significantly accelerated research in tactile representation learning. VT CMC (Yang et al. 2022) models tactile images from GelSight sensors (Yuan, Dong, and Adelson 2017) using contrastive multiview coding (Tian, Krishnan, and Isola 2020). T3 (Zhao et al. 2024a) proposes an architecture with sensor-specific encoders, a shared backbone network, and task-specific decoders, enabling transferable tactile representation learning across multiple sensors. MViTac (Dave, Lygerakis, and Rueckert 2024) uses an InfoNCE loss to jointly optimize visual and tactile features for effective intra- and inter-modal fusion. TLA (Hao et al. 2025) encodes temporal tactile images into composite images and incorporates language reasoning for cross-modal finetuning, thereby promoting generalized tactile-language-action policy learning. SITR (Gupta et al. 2025) combines supervised contrastive learning with physics-based simulation to learn sensor-invariant representations, enabling zero-shot transfer across GelSight sensors. UniTouch (Yang et al. 2024) aligns tactile signals with visual data and introduces learnable sensor-specific tokens to leverage CLIP for multimodal shared representations, supporting various zero-shot tactile tasks. Additionally, UniTouch integrates large language models to facilitate diverse tactile question-answering tasks. TVL-LLaMA (Fu et al. 2024) proposes a tri-modal contrastive-trained tactile encoder aligned with vision and language, and further finetunes LLaMA2 to generate tactile descriptions from visual and tactile inputs. TLV-Link (Cheng et al. 2025) combines teacher-student curriculum learning with contrastive learning for tactile-centric multimodal pretraining. AnyTouch (Feng et al. 2025) integrates static and dynamic information through a hierarchical architecture, incorporating masked modeling, multimodal alignment, and cross-sensor matching for unified multi-sensor tactile representation learning. VIT-LENS-2 (Lei et al. 2024) leverages a pretrained ViT (Dosovitskiy 2020) and modality-specific lens modules for efficient, scalable multimodal tactile learning.
As shown in Table 1, the aforementioned methods fall into three categories: CLIP-Based, LLM-Based, and Custom-Based. The trend shows a shift towards custom end-to-end architectures, but this makes fair benchmarking difficult. Moreover, variations in batch size can significantly affect the stability of multimodal models on tactile tasks (Higuera et al. 2024). Our work focuses on CLIP-Based approaches, which allows us to use a consistent base model (Cherti et al. 2023) and build a standardized RSS evaluation framework for fairer comparisons.
@@ 3 Methodology
We introduce TLV-CoRe, a method for learning collaborative representations that align the tactile (T), visual (V), and language (L) modalities in a shared latent space (see Fig. 2). TLV-CoRe comprises separate encoders for each modality and two modules: a Sensor-Aware Modulator (SAM) in the tactile branch, and a cross-modal Unified Bridging Adapter (UBA) that aligns the modalities. The SAM removes sensor-specific biases (Zhao et al. 2024a; Yang et al. 2024) from tactile features, yielding invariant representations across various tactile sensors. UBA is a lightweight module that projects features from each modality into a common latent space to facilitate alignment. We train the model with symmetric contrastive losses between modalities together with a sensor-invariance loss to enforce consistency across sensors. Finally, we provide theoretical guarantees of TLV-CoRe’s three properties. Detailed proofs can be found in the Appendix. We describe each component in detail below.
@@ 3.1 Tactile Encoder and Sensor-Aware Modulator
The tactile encoder $\mathcal{E}_{T}$ processes raw tactile inputs $x^{T}$ and produces a $d$-dimensional latent representation. Specifically, we implement $\mathcal{E}_{T}$ as a Vision Transformer (ViT) (Dosovitskiy 2020). The tactile image is divided into non-overlapping patches, each patch is linearly embedded, and the resulting sequence is processed through standard transformer blocks. This design enables the encoder to capture spatial patterns and tactile features.
@@
However, tactile images captured under similar conditions can exhibit similar patterns across different sensors (Ou, Chen, and Luo 2024). In such cases, the SAM inadvertently clusters features by sensor identity rather than by underlying tactile content. To address this, we employ tactile-irrelevant decoupled learning that explicitly removes redundant information from the tactile representation.
@@ 3.2 Unified Bridging Adapter for Cross-Modal Alignment
In our implementation we attach exactly L=12L\!=\!12 UBA blocks to each modality so that every sample—regardless of its source—passes through the same number of shared transformations before the final projection. Concretely, the language encoders comprise $12$ transformer layers; we attach one UBA per layer. The vision/tactile branch is deeper (CLIP design (Cherti et al. 2023) with $24$ layers), so we leave the first $12$ layers un‐bridged to preserve low‑level visual/tactile primitives and attach UBAs only to the upper $12$ layers where high‑level semantics emerge.
@@ 3.3 Cross-Modal Contrastive Learning
We train TLV-CoRe with symmetric contrastive losses for each pair of modalities and the sensor-invariance loss. For each pair $(X,Y)\in\{(\mathrm{T},\mathrm{V}),(\mathrm{T},\mathrm{L}),(\mathrm{V},\mathrm{L})\}$, we use a symmetric InfoNCE contrastive loss (Oord, Li, and Vinyals 2018). Given a batch of $N$ aligned triplets $\{(z^{T}_{i},z^{V}_{i},z^{L}_{i})\}_{i=1}^{N}$, the loss for the tactile–vision pair (as an example) is: